# IntGrad MT: Eliciting LLMs’ Machine Translation Capabilities with Sentence Interpolation and Gradual MT

### v2

@@ 7 Conclusion
In this paper, we proposed IntGrad MT, a novel method to enhance the machine translation capabilities of various LLMs. IntGrad MT leverages sentence interpolation to guide models, eliciting stronger translation performance. Experimental results across various models and languages demonstrate that our approach consistently improves translation quality, particularly in low-resource languages, achieving meaningful gains in performance metrics. Our approach is practical in that it does not require extra training and does not conflict with previous methods that utilize other kinds of prompting techniques.
8 Limitations
Despite its success, IntGrad MT introduces some computational overhead, particularly in scenarios involving multiple start sentences or large-scale models. Future work could focus on optimizing computational efficiency or extending the approach to specialized domains and additional low-resource language pairs. Moreover, since sentence interpolation did not perform well in languages other than English, even with GPT-4 (OpenAI et al., 2024), we had to limit our focus to cases where English was the source language. Exploring better prompting techniques to interpolate non-English sentences is a potential future research direction.
Appendix A Few-shot Examples for Sentence Interpolation
In this study, we gave three examples to the interpolation model. We arbitrarily picked start and end sentences from the New York Times news articles.
@@ A.3 Example 3
Appendix B Sample Sentence Interpolation Path
Below is a sample of the interpolation path generated by Qwen2-72b-Intstruct during the experiment.
@@ B.1 Sample 1
Sentence2: “We now have 4-month-old mice that are non-diabetic that used to be diabetic,” he added.
@@
    “We have seen positive transformations in our subjects,” he added.
@@
    “We now have improved conditions in our test subjects,” he added.
@@
    “We now have 4-month-old mice that are non-diabetic that used to be diabetic,” he added.
@@ B.2 Sample 2
Appendix C Prompts and Settings for Translation
ChatGPT
@@ Appendix C Prompts and Settings for Translation
Llama-3.1 70B & 8B
[…]
@@
Mistral-Nemo-Instruct-2407
[…]
@@
Appendix D Settings for Sentence Interpolation
@@
For sentence interpolation Qwen2-72B-Instruct model was used. It was run on one A6000 GPU with 4-bit quantization using transformers library. Temperature and top_p were set to 0.6 and 0.9, respectively.
Appendix E Test Results with Different Metrics
We utilized MetricX (Juraska et al., 2023) and CometKiwi (Rei et al., 2023) to obtain supplementary metrics for evaluation. MetricX and CometKiwi are fine-tuned to predict MQM and DA scores, respectively. Tables 5 and 4 show the results from each model.
Table 4: CometKiwi results of IntGrad MT with the best ablation settings for different target languages.
High Resource Mid Resource Low Resource
MT Model Method DE ZH HI KO SW BN MR
GPT 3.5 0 shot 86.12 85.84 69.12 87.73 83.24 67.93 56.23
3 shot 86.47↑0.34 86.57↑0.72 69.53↑0.41 88.53↑0.79 83.06↓0.18 68.80↑0.86 58.19↑1.96
IntGradDelta 87.43↑1.31 87.61↑1.76 74.33↑5.21 90.92↑3.19 86.15↑2.91 75.60↑7.67 66.21↑9.99
IntGradThr.+Delta 87.15↑1.03 86.35↑0.50 74.14↑5.02 90.66↑2.93 85.98↑2.74 74.94↑7.01 65.60↑9.37
Llama 3.1 70B 0 shot 84.97 - 74.55 - - - -
3 shot 85.38↑0.41 - 75.02↑0.47 - - - -
IntGradDelta 87.15↑2.17 - 78.02↑3.47 - - - -
IntGradThr.+Delta 85.06↑0.09 - 77.93↑3.38 - - - -
Mistral Nemo 0 shot 83.32 82.90 59.72 78.41 42.72 62.03 52.53
3 shot 85.22↑1.89 85.40↑2.49 66.62↑6.90 87.20↑8.79 47.04↑4.32 72.80↑10.78 63.60↑11.07
IntGradDelta 87.09↑3.77 87.13↑4.22 70.61↑10.89 89.35↑10.94 54.46↑11.75 77.91↑15.89 67.89↑15.36
IntGradThr.+Delta 87.00↑3.67 87.13↑4.22 70.22↑10.50 89.32↑10.91 53.39↑10.67 77.76↑15.74 67.77↑15.24
Llama 3.1 8B 0 shot 80.06 - 66.71 - - - -
3 shot 82.11↑2.05 - 69.25↑2.55 - - - -
IntGradDelta 84.71↑4.65 - 73.64↑6.93 - - - -
IntGradThr.+Delta 80.63↑0.56 - 73.24↑6.53 - - - -
Table 5: MetricX results of IntGrad MT with the best ablation settings for different target languages.
High Resource Mid Resource Low Resource
MT Model Method DE ZH HI KO SW BN MR
GPT 3.5 0 shot 0.59 1.10 1.22 0.60 1.32 2.28 2.49
3 shot 0.56↓0.03 1.03↓0.07 1.18↓0.04 0.54↓0.05 1.30↓0.03 2.17↓0.11 2.23↓0.26
IntGradDelta 0.54↓0.04 1.03↓0.07 0.91↓0.31 0.44↓0.16 1.18↓0.15 1.55↓0.73 1.42↓1.07
IntGradThr.+Delta 0.55↓0.04 1.09↓0.02 0.92↓0.30 0.44↓0.16 1.18↓0.14 1.59↓0.69 1.45↓1.04
Llama 3.1 70B 0 shot 0.66 - 0.91 - - - -
3 shot 0.63↓0.02 - 0.80↓0.11 - - - -
IntGradDelta 0.58↓0.08 - 0.68↓0.23 - - - -
IntGradThr.+Delta $0$ 0.68↓0.23 - - - -
Mistral Nemo 0 shot 0.69 1.28 1.66 1.21 7.21 1.91 2.63
3 shot 0.58↓0.11 1.11↓0.16 1.25↓0.42 0.58↓0.63 6.58↓0.63 1.60↓0.31 1.60↓1.03
IntGradDelta 0.54↓0.15 1.09↓0.19 1.11↓0.56 0.54↓0.67 5.45↓1.76 1.13↓0.78 1.26↓1.37
IntGradThr.+Delta 0.55↓0.14 1.09↓0.19 1.13↓0.53 0.54↓0.66 5.56↓1.65 1.13↓0.78 1.26↓1.37
Llama 3.1 8B 0 shot 1.09 - 1.33 - - - -
3 shot 0.81↓0.28 - 1.07↓0.25 - - - -
IntGradDelta 1.13↑0.04 - 0.90↓0.43 - - - -
IntGradThr.+Delta 1.02↓0.07 - 0.92↓0.41 - - - -
Appendix F Samples of interpolation paths
We sampled 100 interpolation paths for En-Ko translation containing more than three sentences and obtained sentence embeddings using SBERT. Each interpolation path was then plotted on a 2D plane using Principal Component Analysis (PCA). Figure 6 shows 24 samples from these 100 paths.
Figure 6: 24 samples of 2D scatter plots. The plots are projected from SBERT embeddings onto a 2D plane using PCA. The X and Y axes of each plot represent the first and second principal components, respectively. Arrows in each plot show the trajectory of sentence shifts from the start sentence (blue-colored dot) to the end sentence (yellow-colored dot).
Appendix G Full Results of Ablation Study

### v3

@@ 8 Conclusion
In this paper, we proposed BridG MT, a novel method to enhance the machine translation capabilities of various LLMs. BridG MT leverages sentence bridging and gradual MT to guide models, eliciting stronger translation performance. Experimental results across various models and languages demonstrate that our approach consistently improves translation quality, particularly in low-resource languages. Our approach is practical in that it does not require extra training and does not conflict with previous methods that utilize other kinds of prompting techniques.
Limitations
BridG MT has a limitation in that it introduces computational overhead due to Sentence Bridging and the recursive nature of Gradual MT. While we have tested various methods to minimize the cost of BridG MT, such as pre-filtering and bridge sampling, future work could focus on further optimizing computational efficiency. Additionally, since the model used for sentence bridging did not perform well in languages other than English, we had to limit our study to cases where English was the source language. Exploring better prompting techniques to interpolate non-English sentences presents a promising direction for future research.
Acknowledgments
This work was supported in part by the National Research Foundation of Korea (NRF) grant (RS-2023-00280883, RS-2023-00222663); by the National Super computing Center with super computing resources including technical support (KSC-2023-CRE-0176, KSC-2024-CRE-0065); by the Korea Institute of Science and Technology Information (KISTI) in 2025 (No.(KISTI) K25L1M1C1), aimed at developing KONI (KISTI Open Neural Intelligence), a large language model specialized in science and technology; and by the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (RS-2025-02263754); by the Korea Medical Device Development Fund grant funded by the Korea government (the Ministry of Science and ICT, the Ministry of Trade, Industry and Energy, the Ministry of Health & Welfare, the Ministry of Food and Drug Safety, 2025의약안003); partially supported by New Faculty Startup Fund from Seoul National University. We thank Deun Lee of Pixel Bridge for her professional assistance in creating the figures used in this paper.
Appendix A Few-shot Examples for Sentence Bridging
In this study, we gave three examples to the bridging model. We arbitrarily picked start and end sentences from the New York Times news articles.
@@ A.3 Example 3
Appendix B Sample Sentence Bridge
Below is a sample of the bridge generated by Qwen2-72b-Intstruct during the experiment.
@@ B.1 Sample 1
Sentence2: “We now have 4-month-old mice that are non-diabetic that used to be diabetic," he added.
@@
    “We have seen positive transformations in our subjects," he added.
@@
    “We now have improved conditions in our test subjects," he added.
@@
    “We now have 4-month-old mice that are non-diabetic that used to be diabetic," he added.
@@ B.2 Sample 2
Appendix C Computational Cost
Tables 7, 8, 9 show the breakdown of the computational cost of BridG MT and the costs of possible scenarios.
Table 7: Breakdown of inference time per component. The statistics are calculated from the FLORES test set into Hindi using Llama-3.1-8b.
Step Average Computation Time per Sentence
Start Sentence Selection (FAISS) 0.09s (for whole dataset)
Pre-Filtering / Post-Filtering (batch size = 8) 6.22s
Sentence Bridging
   w/ Llama-3.2-3B-Instruct 6.03s
   w/ LlaMa-3.1-8B-Instruct 12.94s
   w/ Qwen2-72b-Instruct 26.17s
Translation (Llama-3.1-8B-Instruct)
   Zero-shot 1.74s
   Gradual MT 10.28s
   Gradual MT (w/ bridge sampling) 2.29s
   3-shot (for the final output) 2.18s
Baselines
   0 shot 1.74s
   30 shot 2.33s
   50 shot 2.65s
Table 8: Comparison of sentence bridging models in terms of inference time and average xCOMET score across different MT models and languages.
Bridging Model Time for Sentence Bridging Time for Overall Process Average Score
Qwen2-72B-Instruct 26.17s 44.41s 81.37
Llama-3-1-8B-Instruct 12.94s (49.45%) 31.18s (70.21%) 80.40
Llama-3-2-3B-Instruct 6.03s (23.04%) 24.27s (54.65%) 80.00
0-shot Baseline 1.74s 1.74s 76.35
50-shot Baseline 2.65s 2.65s 78.71
Table 9: Comparison of using the full bridge and bridge sampling in terms of inference time and xCOMET score across different MT models and languages.
Bridge Sampling Time for Gradual MT Time for Overall Process Average Score (xCOMET)
Full Path 3 start sentence 10.28s 44.41s 81.37
Bridge Sampling 1 start sentence 2.29s (22.28%) 36.42s (82.01%) 79.96
0-shot Baseline 1.74s 76.35
50-shot Baseline 2.65s 78.71
Table 10: Comparison of filtering strategies in terms of time and xCOMET score across different MT models and languages.
Filtering Time for Overall Process Average Score (xComet)
post-filtering 46.59s 81.37
pre-filtering 33.46s 80.25
0-shot Baseline 1.74s 76.35
50-shot Baseline 2.65s 78.71
Appendix D Prompts and Settings for Translation
D.1 ChatGPT
@@
D.2 Llama-3.1 70B & 8B
Llama-3.1 Instruct models were run on one A6000 GPU, using transformers library. 70B model were 4-bit quantized.Temperature and top_p were set to 0.6 and 0.9, respectively. The actual prompt is as follows:
@@
D.3 Mistral-Nemo-Instruct-2407
Mistral-Nemo-Instruct-2407 was run on one A6000 GPU, using transformers library. Temperature and top_p were set to 0.6 and 0.9, respectively. The actual prompt is as follows:
User: You will be provided with a sentence in English, and your task is to translate it into $\langle$ Target Language $\rangle$.
@@ D.3 Mistral-Nemo-Instruct-2407
Appendix E Settings for Sentence Bridging
@@
Qwen2-72B-Instruct model was used for main experiments. It was run on one A6000 GPU with 4-bit quantization using transformers library. Temperature and top_p were set to 0.6 and 0.9, respectively.
Llama-3.1-8B-Instruct model was used for additional experiments. It was run on one A6000 GPU using transformers library. Temperature and top_p were set to 0.6 and 0.9, respectively.
Appendix F Full Results with Different Metrics
Table 11, 12, 13, 14, 15, 16 shows full results of our main experiment with different metrics; xCOMET, CometKiwi, MetricX, BLEURT, BLEU and ChrF.
Table 11: Full xCOMET results across different sentence bridging models, translation models and target languages.
GPT-3.5-turbo-0125 Llama-3.1-70B Llama-3.1-8B Mistral-Nemo
SETTING DE ZH HI KO SW BN MR DE HI DE HI DE ZH HI KO SW BN MR
BASELINE 0 Shot 97.63 91.30 71.89 89.48 81.23 68.73 44.53 97.33 79.41 94.99 69.93 96.70 88.01 66.78 81.43 38.97 71.89 43.99
BASELINE 15 Shot 98.01 92.16 73.13 90.73 81.59 69.70 45.54 97.10 77.95 96.26 73.70 97.68 90.94 69.54 88.71 42.79 73.60 52.26
BASELINE 30 Shot 97.92 91.87 72.98 90.48 81.62 68.58 45.33 96.73 74.67 96.55 73.74 97.72 90.79 69.84 88.06 42.59 74.10 52.52
BASELINE 50 Shot 97.99 91.95 72.85 90.93 82.10 67.72 44.84 96.56 72.68 96.42 73.36 97.72 90.92 69.35 88.29 42.53 73.80 52.04
BridG MT w/ Qwen2-72b
Post Filtering 98.04 92.63 77.90 92.57 83.65 75.41 53.11 97.93 84.45 97.25 78.30 97.70 90.99 73.95 89.45 45.67 78.35 57.24
Pre & Post Filtering 97.86 92.35 77.63 92.07 83.34 75.14 51.51 97.73 84.31 96.99 77.37 97.70 90.99 73.82 89.34 45.18 77.83 54.55
BridG MT w/ Llama-3.1-8b
Post Filtering 97.97 92.38 77.24 91.85 83.62 74.24 52.27 97.87 83.97 96.93 77.83 97.52 90.40 72.34 87.83 43.33 76.73 52.93
Pre & Post Filtering 97.95 91.54 76.96 91.42 83.52 74.06 50.86 97.73 83.90 96.82 73.09 97.41 90.40 71.19 87.42 43.09 76.23 51.29
BridG MT w/ Llama-3.2-3b
Post Filtering 97.97 92.35 76.44 91.48 83.49 74.35 52.13 97.85 83.65 97.14 77.65 97.42 90.04 71.62 86.55 42.30 75.83 51.66
Pre & Post Filtering 97.92 92.29 76.39 91.21 82.74 74.10 52.04 97.61 83.28 96.86 76.84 97.42 90.04 71.46 86.33 42.14 75.75 51.05
QA Decode 97.45 91.63 75.45 90.19 78.26 73.90 45.82 97.37 82.32 88.08 77.67 96.33 87.18 64.42 80.34 28.86 68.59 35.37
MAPS 97.68 92.07 - - - - - 96.53 - 79.10 - 85.75 88.71 - - - - -
Table 12: Full CometKiwi results across different sentence bridging models, translation models and target languages.
GPT-3.5-turbo-0125 Llama-3.1-70B Llama-3.1-8B Mistral-Nemo
SETTING DE ZH HI KO SW BN MR DE HI DE HI DE ZH HI KO SW BN MR
BASELINE 0 Shot 86.12 85.84 69.12 87.73 83.24 67.93 56.23 84.97 74.55 80.06 66.71 83.32 82.90 59.72 78.41 42.72 62.03 52.53
BASELINE 15 Shot 86.71 86.71 69.72 88.85 82.99 69.65 58.37 84.11 72.94 81.64 69.58 85.46 85.73 66.77 87.48 47.51 72.81 63.73
BASELINE 30 Shot 86.44 86.74 69.51 88.86 83.12 68.39 57.71 83.22 70.19 82.17 69.40 85.38 85.74 67.00 86.88 48.08 73.00 63.35
BASELINE 50 Shot 86.61 86.81 69.30 89.18 83.42 67.98 57.32 82.70 68.28 81.95 69.05 85.30 86.04 66.78 87.08 47.61 72.83 63.09
BridG MT w/ Qwen2-72b
Post Filtering 87.42 87.64 74.30 90.88 86.09 75.34 66.01 87.25 78.04 85.40 74.00 86.63 86.89 70.96 89.51 54.71 77.65 68.16
Pre & Post Filtering 86.80 87.23 74.12 90.65 85.92 75.09 64.94 86.33 77.96 84.44 73.52 86.63 86.89 70.85 89.48 54.40 77.26 66.82
BridG MT w/ Llama-3.1-8b
Post Filtering 87.28 87.42 73.83 90.36 85.76 74.30 64.86 87.16 77.91 84.86 73.25 86.48 86.39 69.42 88.17 51.82 75.79 64.70
Pre & Post Filtering 87.20 86.27 73.66 90.12 85.73 74.07 64.01 86.28 77.82 84.53 70.18 86.11 86.39 68.75 87.98 51.49 75.42 63.82
BridG MT w/ Llama-3.2-3b
Post Filtering 87.13 87.38 73.24 90.29 85.58 74.42 64.74 87.19 77.94 85.31 73.30 86.20 85.97 68.62 87.63 50.19 74.83 6371
Pre & Post Filtering 86.91 87.26 73.21 90.08 84.95 74.22 64.67 85.98 77.64 84.41 72.84 86.20 85.97 68.54 87.47 50.00 7480 6349
QA Decode 85.07 85.84 68.52 88.16 82.98 72.29 63.61 84.85 73.31 82.52 69.60 81.93 82.20 55.94 77.56 39.83 60.05 50.49
MAPS 85.61 86.39 - - - - - 83.48 - 71.19 - 71.06 84.61 - - - - -
Table 13: Full BLEURT results across different sentence bridging models, translation models and target languages.
GPT-3.5-turbo-0125 Llama-3.1-70B Llama-3.1-8B Mistral-Nemo
SETTING DE ZH HI KO SW BN MR DE HI DE HI DE ZH HI KO SW BN MR
BASELINE 0 Shot 78.58 73.70 68.31 68.77 75.80 67.95 68.03 77.57 71.34 74.69 66.96 76.27 69.87 60.95 59.13 51.51 62.29 64.20
BASELINE 15 Shot 79.27 74.16 68.49 69.41 75.84 68.46 69.29 77.12 70.61 75.69 68.54 77.95 72.05 66.52 66.94 54.71 70.14 70.75
BASELINE 30 Shot 79.06 74.19 68.75 69.41 76.01 67.87 69.00 76.35 69.32 75.99 68.62 78.07 72.30 66.64 66.56 55.00 69.79 70.91
BASELINE 50 Shot 79.10 74.20 68.74 69.54 76.13 67.82 68.56 76.10 68.68 75.80 68.41 77.89 72.40 66.46 66.49 55.19 70.15 70.46
BridG MT w/ Qwen2-72b
Post Filtering 79.11 74.11 69.91 70.28 76.51 70.85 71.22 78.56 72.39 76.98 69.20 77.91 71.38 66.85 67.17 56.50 71.71 71.40
Pre & Post Filtering 78.88 73.99 69.80 70.05 76.40 70.70 70.84 78.17 72.35 76.49 68.92 77.91 71.38 66.82 67.15 56.39 71.41 70.66
BridG MT w/ Llama-3.1-8b
Post Filtering 79.13 74.19 69.75 69.99 76.31 70.54 70.65 78.61 72.28 76.75 68.88 77.67 71.04 65.94 66.00 55.16 70.71 70.04
Pre & Post Filtering 79.11 73.85 69.69 69.59 76.24 70.41 70.32 78.19 72.33 76.61 67.85 77.48 71.04 65.58 65.79 55.14 70.42 69.56
BridG MT w/ Llama-3.2-3b
Post Filtering 78.94 74.16 69.51 69.62 76.31 70.10 70.61 78.55 72.30 76.93 68.84 77.65 70.94 65.59 65.03 54.61 70.02 69.45
Pre & Post Filtering 78.89 74.10 69.51 69.52 76.08 70.01 70.59 78.07 72.28 76.42 68.69 77.65 70.94 65.53 6492 54.56 70.00 69.33
QA Decode 78.72 74.07 69.83 69.45 76.60 71.63 72.04 78.41 72.67 76.78 70.00 76.11 69.81 60.85 59.18 51.42 61.93 64.32
MAPS 79.08 74.26 - - - - - 77.40 - 70.81 - 63.18 71.69 - - - - -
Table 14: Full MetricX results across different sentence bridging models, translation models and target languages.
GPT-3.5-turbo-0125 Llama-3.1-70B Llama-3.1-8B Mistral-Nemo
SETTING DE ZH HI KO SW BN MR DE HI DE HI DE ZH HI KO SW BN MR
BASELINE 0 Shot 0.5864 1.1045 1.2189 0.5998 1.3240 2.2832 2.4860 0.6576 0.9091 1.0922 1.3274 0.6904 1.2751 1.6646 1.2064 7.2095 1.9082 2.6305
BASELINE 15 Shot 0.5428 1.0382 1.1840 0.5156 1.3116 2.0955 2.1551 0.7084 0.9696 0.7956 1.1012 0.5839 1.0667 1.2366 0.5462 6.3982 1.6427 1.6782
BASELINE 30 Shot 0.5477 1.0515 1.1629 0.5000 1.3188 2.1823 2.2366 0.7742 1.1978 0.7188 1.0863 0.5884 1.0718 1.2482 0.5842 6.3487 1.6230 1.8343
BASELINE 50 Shot 0.5532 1.0531 1.1319 0.5038 1.2999 2.2167 2.3578 0.7932 1.2621 0.7150 1.0899 0.5820 1.0725 1.2728 0.5568 6.5473 1.6864 1.8806
BridG MT w/ Qwen2-72b
Post Filtering 0.5432 1.0154 0.9179 0.4343 1.1661 1.5156 1.4508 0.5701 0.6854 0.6825 0.8481 0.5639 1.0808 1.0749 0.5559 5.3817 1.1181 1.2351
Pre & Post Filtering 0.5609 1.0370 0.9264 0.4460 1.1716 1.5285 1.5104 0.6058 0.6895 0.7356 0.8846 0.5639 1.0808 1.0809 0.5606 5.4107 1.1342 1.2881
BridG MT w/ Llama-3.1-8b
Post Filtering 0.5494 1.0534 0.9157 0.4822 1.1922 1.6673 1.5956 0.5868 0.7184 0.7214 0.9050 0.5784 1.1299 1.2340 0.6131 5.8600 1.3074 1.5470
Pre & Post Filtering 0.5497 1.0896 0.9205 0.4948 1.1984 1.6793 1.6469 0.6022 0.7205 0.7467 1.0581 0.6008 1.1299 1.2725 0.6371 5.8815 1.3304 1.5953
BridG MT w/ Llama-3.2-3b
Post Filtering 0.5471 1.0415 0.9824 0.4673 1.2055 1.6778 1.6031 0.5681 0.7210 0.6756 0.8896 0.6017 1.1462 1.2412 0.7434 6.1683 1.3213 1.6552
Pre & Post Filtering 0.5532 1.0404 0.9827 0.4759 1.2334 1.6827 1.6037 0.6146 0.7326 0.7168 0.9233 0.6017 1.1462 1.2521 0.7620 6.1826 1.3214 1.6684
QA Decode 0.5764 1.0660 0.9071 0.4992 1.2059 1.3878 1.2131 0.5948 0.7066 0.7348 0.7422 0.6983 1.2705 1.6804 1.1841 7.2072 1.9488 2.5831
MAPS 0.5516 1.0451 - - - - - 0.6939 - 5.9713 - 3.6419 1.1743 - - - - -
Table 15: Full BLEU results across different sentence bridging models, translation models and target languages.
GPT-3.5-turbo-0125 Llama-3.1-70B Llama-3.1-8B Mistral-Nemo
SETTING DE ZH HI KO SW BN MR DE HI DE HI DE ZH HI KO SW BN MR
BASELINE 0 Shot 40.45 45.58 23.06 27.86 32.93 9.99 5.94 38.71 29.09 30.81 21.54 35.70 38.83 17.46 20.17 12.13 8.35 5.31
BASELINE 15 Shot 40.86 46.02 22.98 28.81 33.80 9.67 7.54 37.33 25.54 32.13 22.14 36.89 40.65 20.88 25.08 14.40 11.03 7.30
BASELINE 30 Shot 40.86 46.00 23.08 29.08 33.57 9.10 6.94 36.13 23.29 32.75 22.36 36.96 41.03 21.05 25.45 14.44 11.13 7.45
BASELINE 50 Shot 40.96 45.98 23.51 28.90 33.95 9.75 6.08 35.39 22.84 32.61 22.37 36.72 41.14 21.05 25.28 13.75 11.16 7.69
BridG MT w/ Qwen2-72b
Post Filtering 40.87 46.35 23.72 29.01 33.30 11.06 8.00 38.42 28.08 32.43 22.12 35.79 37.76 18.53 23.96 14.09 10.06 7.19
Pre & Post Filtering 40.65 46.05 23.60 28.65 33.18 10.94 7.96 38.68 28.13 32.29 22.19 35.79 37.76 18.52 23.92 13.96 10.13 6.85
BridG MT w/ Llama-3.1-8b
Post Filtering 40.86 46.27 23.97 28.87 33.57 10.93 7.94 39.11 28.11 32.23 22.09 35.65 37.77 18.71 23.57 13.39 10.11 6.80
Pre & Post Filtering 40.82 45.65 23.87 28.22 33.46 10.81 7.79 39.08 28.25 32.15 21.97 36.23 37.77 18.51 23.83 13.35 10.07 6.47
BridG MT w/ Llama-3.2-3b
Post Filtering 40.79 46.18 23.53 28.66 33.30 10.86 8.42 38.95 28.36 32.37 22.09 36.08 37.49 18.82 22.89 12.96 9.87 6.07
Pre & Post Filtering 40.68 46.01 23.55 28.33 33.19 10.70 8.47 38.79 28.43 32.07 22.20 36.08 37.49 18.73 23.07 12.87 9.84 6.10
QA Decode 41.07 46.02 23.65 28.54 33.59 11.16 8.57 39.56 30.06 34.51 23.66 35.81 38.67 17.46 20.39 12.11 7.98 5.40
MAPS 40.83 46.15 - - - - - 38.45 - 11.04 - 17.74 40.63 - - - - -
Table 16: Full ChrF results across different sentence bridging models, translation models and target languages.
GPT-3.5-turbo-0125 Llama-3.1-70B Llama-3.1-8B Mistral-Nemo
SETTING DE ZH HI KO SW BN MR DE HI DE HI DE ZH HI KO SW BN MR
BASELINE 0 Shot 66.91 39.28 50.42 34.72 62.20 42.82 37.01 65.77 55.83 60.95 48.38 63.49 33.81 42.18 27.60 41.66 38.18 34.12
BASELINE 15 Shot 67.16 39.90 50.15 35.58 62.88 42.71 39.71 64.91 52.34 61.64 49.09 64.26 35.68 46.90 32.24 43.19 44.58 39.71
BASELINE 30 Shot 67.15 39.82 50.71 35.66 62.67 42.09 38.62 64.25 50.53 61.73 49.00 64.31 36.25 47.16 32.39 43.27 44.67 39.77
BASELINE 50 Shot 67.05 39.78 50.60 35.76 62.80 42.41 37.68 63.73 50.00 61.81 49.12 64.01 36.26 47.13 32.34 42.94 44.50 39.67
BridG MT w/ Qwen2-72b
Post Filtering 67.35 39.76 51.40 35.85 62.65 44.52 40.70 65.72 55.27 61.33 49.45 63.48 32.96 45.63 31.26 43.71 44.01 39.52
Pre & Post Filtering 67.11 39.51 51.31 35.48 62.56 44.42 40.48 65.76 55.28 61.52 49.30 63.48 32.96 45.61 31.26 43.68 43.96 39.04
BridG MT w/ Llama-3.1-8b
Post Filtering 67.24 39.69 51.53 35.61 62.62 44.31 40.34 66.16 55.26 61.63 49.39 63.49 32.90 45.24 30.70 42.99 43.52 38.10
Pre & Post Filtering 67.20 39.31 51.44 35.04 62.56 44.24 40.18 66.01 55.32 61.61 48.99 63.81 32.90 45.05 30.97 43.05 43.52 37.87
BridG MT w/ Llama-3.2-3b
Post Filtering 67.18 39.64 51.32 35.37 62.50 44.20 40.83 66.06 55.44 61.54 49.33 63.65 32.83 45.05 30.33 42.87 43.05 38.10
Pre & Post Filtering 67.17 39.47 51.33 35.09 62.44 44.13 40.82 65.92 55.57 61.45 49.34 63.65 32.83 44.94 30.51 42.88 43.04 38.06
QA Decode 67.26 39.65 51.19 35.31 62.68 44.28 41.06 66.47 56.82 62.92 50.60 63.47 33.81 41.99 27.75 41.76 38.07 34.05
MAPS 67.29 39.69 - - - - - 65.17 - 47.66 - 54.36 34.68 - - - - -
Figure 4: MetricX scores for few-shot translations (bars) and BridG MT translations (horizontal line). Lower score indicates better performance. Both BridG MT and few-shot scores represent the average MetricX scores of GPT-3.5 and Mistral-Nemo.
Appendix G Results from Ablation